# Data-driven decision-making under uncertainty with entropic risk measure

### v1

@@ 5 Distributionally Robust Optimization
where $\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij}):=\mathop{\rm inf}_{{{\boldsymbol{\xi}}}}{{\boldsymbol{\varphi}}}_{ij}^{\top}{{\boldsymbol{\xi}}}-\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ is the partial concave conjugate of $\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$.
In Appendix A, we provide reformulations of distributionally robust newsvendor and regression problems as exponential cone programs. However, robustifying the SAA problem does not necessarily change the optimal decision, as highlighted in the following remark:
Remark 1
By translation invariance of entropic risk measure, minimizing the worst-case entropic risk of loss function that is separable in ${{\boldsymbol{\xi}}}$ and ${{\boldsymbol{z}}}$ is equivalent to solving the SAA problem.
Furthermore, by assuming that the tails of loss $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$ are exponentially bounded as in Assumption 3 with $C>p$, we can ensure that the rate of convergence of risk of SAA problem to the true risk is $\text{O}(N^{-\frac{p}{2}})$ for sufficiently large $N$ (see Lemma 2 in the Appendix B). This also ensures the convergence of $\rho_{\mbox{\scriptsize DRO}}$ to $\rho^{*}$ in probability (see Theorem 5 in Appendix B).
@@
1:function KfoldCV($K,\mathcal{D}_{N},\epsilon$)
2:     $\mathcal{X}\leftarrow\emptyset$
3:     for $f\leftarrow 1$ to $K$ do
4:         $\mathcal{D}_{-f}\leftarrow\mathcal{D}_{N}\setminus\mathcal{D}_{f}$triangleright Training data (all samples except those in fold $f$)
5:         Solve problem (18) with scenarios in $\mathcal{D}_{-f}$ and radius $\epsilon$ to get ${{\boldsymbol{z}}}_{-f}$
6:         $\mathcal{X}\leftarrow\mathcal{X}\cup\{\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}}_{-f})\mid{{\boldsymbol{\xi}}}\in\mathcal{D}_{f}\}$
7:     end for
8:     return $\mathcal{X}$
9:end function
5.1 Calibrating the radius $\epsilon$
In this section, we describe the procedure for calibrating the radius parameter $\epsilon$. Let $\mathcal{G}$ be a grid of potential radius values, including 00. For each radius $\epsilon\in\mathcal{G}$, Algorithm 4 applies the K-fold cross-validation algorithm to generate the validation scenarios $\mathcal{X}$. Specifically, we partition the dataset $\mathcal{D}_{N}$ into $K$ folds, where each fold is denoted by $f\in\{1,2,\cdots,K\}$.
The model is trained on all but one fold, $(\mathcal{D}_{-f})$, and the loss is evaluated on the remaining fold, $(\mathcal{D}_{f})$, to produce validation losses. To this end, we solve (18) over dataset $\mathcal{D}_{-f}$ to find the optimal decision ${{\boldsymbol{z}}}_{-f}$. The loss of decision ${{\boldsymbol{z}}}_{-f}$ is then computed on each sample, ${{\boldsymbol{\xi}}}\in\mathcal{D}_{f}$, in the validation fold $f$. This process is repeated for all folds and the resulting loss scenarios are aggregated to form a set $\mathcal{X}$ of size equal to training sample size $N$.
Algorithm 5 calculates the entropic risk $\rho_{\mathcal{X}}$ across the scenarios in $\mathcal{X}$ for each radius $\epsilon\in\mathcal{G}$. A GMM is then fitted to the scenarios in $\mathcal{X}$, and and bootstrapping is used to estimate the bias, which is added to $\rho_{\mathcal{X}}$ to obtain the bias-corrected risk $\hat{\rho}(\epsilon)$. This procedure is repeated over all radius $\epsilon\in\mathcal{G}$ and the optimal radius $\epsilon^{*}$ is chosen as the one minimizing $\hat{\rho}(\epsilon)$. The GMM can be fitted using MLE, entropic risk matching (Match), or by matching the tails (EVT).
In contrast, traditional CV sets $\delta_{\epsilon}$ to 00 for all $\epsilon\in\mathcal{G}$. By Jensen’s inequality, this underestimates the true risk in the validation set, potentially leading to errors in radius selection. In Example 3, the projects can be seen as corresponding to different regularization parameters, with lower values of $\epsilon$ representing riskier projects. Traditional CV would select project 3, whereas our approach, based on bias-corrected estimators, selects project 1.
@@ 5.1 Calibrating the radius $\epsilon$
1:function Radius Tuning(${\mathbb{P}},K,N$)
2:     $\mathcal{D}_{N}\leftarrow\text{generate }N\text{ i.i.d.~{}samples from }{\mathbb{P}}$
3:     for $\epsilon\in\mathcal{G}$ do
[…]
5:         $\rho_{\mathcal{X}}\leftarrow$ compute entropic risk of loss scenarios in $\mathcal{X}$
6:         $\delta_{\epsilon}\leftarrow\text{Bootstrapping}(\mathcal{X})$
7:         $\hat{\rho}(\epsilon)\leftarrow\rho_{\mathcal{X}}+\delta_{\epsilon}$
8:     end for
9:     $\epsilon^{*}\leftarrow\arg\mathop{\rm min}_{\epsilon\in\mathcal{G}}\hat{\rho}(\epsilon)$
10:end function
6 Distributionally robust insurance pricing
The International Panel on Climate Change (IPCC) advocates using financial instruments like catastrophe insurance to mitigate risks associated with rare, high-impact events such as floods, earthquakes, and wildfires. Such instruments are crucial for climate change adaptation (Linnerooth-Bayer and Hochrainer-Stigler,, 2015). These events have become more frequent due to climate change, making reliable risk estimation critical. However, the rarity of such events leads to limited data, which poses challenges for accurate risk assessment. High insurance premiums often deter homeowners from purchasing insurance, with demand typically spiking only after catastrophic events (Gallagher,, 2014). The correlated risks of floods, earthquakes, and wildfires can lead to significant payouts for insurers, limiting the availability of flood insurance in commercial markets. For instance, the US National Flood Insurance Program (NFIP) faces over \$20 billion deficit (Marcoux and H Wagner,, 2023) due to underpriced premiums and low uptake, highlighting the need for optimal pricing that accounts for homeowners’ behavioral responses (Kousky and Cooke,, 2012).
We consider a model of insurance pricing with one risk-averse insurer and $M$ representative risk-averse homeowners, based on the framework introduced by Bernard et al., (2020). In their model, both the insurer and homeowners are assumed to be expected utility maximizers, with loss distributions and risk aversion parameters known to the insurer. We depart from these assumptions by only providing samples of the loss distribution to both the insurer and homeowners.
Let $\alpha_{i}$ be the risk aversion of homeowner $i$, and let the insurer’s risk aversion parameter be $\alpha_{0}$. The uncertain loss faced by homeowner $i$ is represented by $\xi_{i}$. The insurer offers a policy $(I_{i},\pi_{i})$ to homeowner $i$ where the indemnity function $I_{i}:=z_{i}\xi_{i}$ specifies the coverage provided, and $\pi_{i}$ is the premium paid by the homeowner.
If homeowner $i$ rejects the insurance policy, the entropic risk is $\rho_{{\mathbb{P}}_{i}}\left(\xi\right)$.
If homeowner $i$ accepts the insurance policy, the entropic risk is $\rho_{{\mathbb{P}}_{i}}\left(\xi_{i}-I(\xi_{i})+\pi_{i}\right)$.
Let ${\mathbb{P}}_{i}$ denote each homeowner’s loss distribution $\xi_{i}$. The insurer’s demand response model assumes that homeowner $i$ will accept the insurance policy if the empirical entropic risk with insurance is less than the entropic risk without insurance:
$\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi_{i}-I(\xi_{i})+\pi_{i}\right)\leq\rho_{\hat{{\mathbb{P}}}_{i,N}}\left(\xi\right),$
where ^$\hat{{\mathbb{P}}}_{i,N}$ is the empirical distribution of the losses faced by homeowner $i$. This model assumes that the decision to accept the insurance policy is based solely on the empirical distribution of the losses. Alternative demand response models could be considered, allowing for a corresponding analysis to be conducted.
The entropic risk of the insurer is given by:
$\rho_{{\mathbb{P}}}\left(\sum_{i=1}^{N}z_{i}\xi_{i}-\sum_{i=1}^{N}\pi_{i}\right).$

### v3

@@ 5 Distributionally Robust Optimization
where $\ell_{j*}({{\boldsymbol{z}}},{{\boldsymbol{\varphi}}}_{ij}):=\mathop{\rm inf}_{{{\boldsymbol{\xi}}}}\{{{\boldsymbol{\varphi}}}_{ij}^{\top}{{\boldsymbol{\xi}}}-\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})\}$ is the partial concave conjugate of $\ell_{j}({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})$, and $\|\cdot\|_{*}$ denotes the dual norm.
The DRO reformulation and reformulation technique simplify significantly when the loss function is either piecewise linear or linear, rather than piecewise convex in $\boldsymbol{z}$ and concave in $\boldsymbol{\xi}$. The following corollary presents these special cases.
Corollary 6
Let $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}):=\mathop{\rm max}_{k\in\mathcal{K}}\left\{a_{k}({{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})+b_{k}\right\}$ be a piecewise linear function for given parameters $a_{k}$ and $b_{k}$. Then, the DRO problem (12) with type-$\infty$ Wasserstein ambiguity set is equivalent to
$\rho_{\texttt{DRO}}:=\quad\mathop{\rm min}$ $\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}t_{i}\right)$ (14)
s.t. $\boldsymbol{t}\in\mathbb{R}^{N},\,{{\boldsymbol{z}}}\in\mathcal{Z}$
$\exp\left(\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})+b_{k}\right)+\epsilon\|a_{k}{{\boldsymbol{z}}}\|_{*}\right)\leq t_{i}$ $\forall i,k\in\mathcal{K}$
which for a linear loss $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}})={{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}$ can be further simplified to
$\mathop{\rm min}_{{{\boldsymbol{z}}}\in\mathcal{Z}}\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\exp(\alpha{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})\right)+\epsilon\left\lVert{{\boldsymbol{z}}}\right\rVert_{*}.$ (15)
Proof. Here, we provide an alternative proof for the piecewise linear loss functions $\ell({{\boldsymbol{z}}},{{\boldsymbol{\xi}}}):=\mathop{\rm max}_{k\in\mathcal{K}}\left\{a_{k}({{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})+b_{k}\right\}$ that does not rely on Fenchel duality (Ben-Tal et al.,, 2015). The supremum of $\exp(\mathop{\rm max}_{k\in\mathcal{K}}\left\{\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})+b_{k}\right)\right\})$ over the set $\{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon\}$ is given by:
$\mathop{\rm sup}_{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}$ $\exp\left(\mathop{\rm max}_{k\in\mathcal{K}}\left\{\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})+b_{k}\right)\right\}\right)$
$=\mathop{\rm sup}_{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}-\hat{{{\boldsymbol{\xi}}}}_{i}\|\leq\epsilon}\mathop{\rm max}_{k\in\mathcal{K}}\left\{\exp\left(\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}})+b_{k}\right)\right)\right\}$
$=\mathop{\rm max}_{k\in\mathcal{K}}\left\{\exp\left(\alpha\left(a_{k}{{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i}+b_{k}\right)+\alpha\mathop{\rm sup}_{{{\boldsymbol{\xi}}}:\|{{\boldsymbol{\xi}}}\|\leq\epsilon}\left(a_{k}{{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}\right)\right)\right\}$
$=\mathop{\rm max}_{k\in\mathcal{K}}\left\{\exp\left(\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})+b_{k}\right)+\alpha\epsilon\|a_{k}{{\boldsymbol{z}}}\|_{*}\right)\right\},$
[…]
$\frac{1}{\alpha}\log\left(\frac{1}{N}\sum_{i=1}^{N}\mathop{\rm max}_{k\in\mathcal{K}}\left\{\exp\left(\alpha\left(a_{k}({{\boldsymbol{z}}}^{\top}\hat{{{\boldsymbol{\xi}}}}_{i})+b_{k}\right)+\alpha\epsilon\|a_{k}{{\boldsymbol{z}}}\|_{*}\right)\right\}\right).$
So, with a piecewise linear loss function, problem (12) is equivalent to the convex optimization problem in (14). Further, specializing the result to a linear loss function $\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}})={{\boldsymbol{z}}}^{\top}{{\boldsymbol{\xi}}}$, problem (12) is equivalent to the regularized risk-averse SAA problem in (15).
It is interesting to see that for the linear case, the DRO problem reduces to the regularized SAA problem where the regularization penalty is controlled by the size $\epsilon$ of the ambiguity set and that the type of penalty depends on the dual of the norm used to define the ambiguity set. To complement these results, Appendix A.8 provides reformulations of the distributionally robust newsvendor and regression problems as exponential cone programs.
Our next theorem formalizes that as the sample size $N$ tends to infinity, the DRO value $\rho_{\texttt{DRO}}$ with a properly chosen radius will converge to the true optimal risk $\rho^{*}$ in probability. The proof follows from showing that for Lipschitz continuous (in ${{\boldsymbol{z}}}$) loss functions, $\rho_{\texttt{SAA}}\leq\rho_{\texttt{DRO}}\leq\rho_{\text{{SAA}}}+L\epsilon$ and using Proposition 3 that establishes that $\rho_{\texttt{SAA}}$ converges to $\rho^{*}$ at the rate $\mathcal{O}(1/\sqrt{N})$ for locally Lipschitz continuous (in ${{\boldsymbol{\xi}}}$) loss functions. Finally, choosing the radius to decay at the rate $\mathcal{O}(1/\sqrt{N})$ preserves the rate of convergence of SAA.
Theorem 7
Suppose that Assumption 3 holds. Then for any $\gamma>0$ and using $\mathcal{B}_{\infty}(c/\sqrt{N})$, for some $c>0$, then there exists a constant $A>0$ such that
${\mathbb{P}}\left(\lvert\rho_{\texttt{DRO}}-\rho^{*}\rvert\geq\frac{A}{\sqrt{N\gamma}\alpha\exp(\alpha\rho^{*})}+\frac{c}{\sqrt{N}}\right)\leq\gamma,$
as long as $N$ is sufficiently large. Consequently, $\rho_{\texttt{DRO}}\rightarrow\rho^{*}$ in probability.
While Theorem 7 provides a rate for $\epsilon$ that ensures convergence of the DRO risk to the true optimal risk in probability, the values of the constants depend on the unknown underlying probability distribution. In practice, $\epsilon$ needs to be estimated using CV. However, as discussed in the previous sections, estimating true risk from finite data is challenging. To address this, we next employ the bias-aware estimation procedure described in Section 4.
5.1 Bias-aware cross validation
A common approach to select the radius $\epsilon$ of the ambiguity set is through $K$-fold CV. For each $\epsilon\in\mathcal{E}$, $K$-fold CV aims to estimate the true performance of policy ${{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{N},\epsilon)$ resulting from problem (13), i.e., $\rho_{\mathbb{P}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{N},\epsilon),{{\boldsymbol{\xi}}}))$, and subsequently select the $\epsilon$ that minimizes this risk. The approach divides the dataset into $K$ folds. For each fold, we optimize the DRO model on $K-1$ folds and evaluate the solution’s performance on the remaining fold, repeating this process for all folds. Specifically, for each candidate value of $\epsilon$, the model in problem (13) is solved using the training data ^$\hat{\mathbb{P}}_{-k}^{K}$ from all folds except the $k$-th fold to determine ${{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon)$ which is then evaluated on the validation data to obtain $\rho_{{{\boldsymbol{\xi}}}\sim\hat{\mathbb{P}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon),{{\boldsymbol{\xi}}}))$, where ^$\hat{\mathbb{P}}_{k}^{K}$ denotes the empirical distribution of scenarios in fold $k$. The resulting estimator for a given radius $\epsilon$ is then given $\rho_{k\sim U(K)}(\rho_{{{\boldsymbol{\xi}}}\sim\hat{\mathbb{P}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon),{{\boldsymbol{\xi}}})))$, where $U(K)$ is the uniform distribution over the set $\{1,2,\ldots,K\}$. Since the goal is to minimize risk, we choose $\epsilon$ that minimizes the validation risk, i.e., $\epsilon^{*}=\mathop{\rm argmin}_{\epsilon\in\mathcal{E}}\rho_{k\sim U(K)}(\rho_{{{\boldsymbol{\xi}}}\sim\hat{\mathbb{P}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon),{{\boldsymbol{\xi}}})))$. However, for each value of radius $\epsilon$ and choice of $K$, the following proposition shows that the entropic risk estimator $\rho_{k\sim U(K)}(\rho_{{{\boldsymbol{\xi}}}\sim\hat{\mathbb{P}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon),{{\boldsymbol{\xi}}})))$ based on the K-fold CV, underestimates the entropic risk of the policy constructed using $N(1-\frac{1}{K})$ data points.
Proposition 8
Given $\epsilon\in\mathcal{E}$,
${\mathbb{E}}[\rho_{k\sim U(K)}(\rho_{{{\boldsymbol{\xi}}}\sim\hat{\mathbb{P}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{-k}^{K},\epsilon),{{\boldsymbol{\xi}}})))]<\rho_{\mathbb{P}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{N(1-\frac{1}{K})},\epsilon),{{\boldsymbol{\xi}}})).$ (16)
The proof of the above proposition follows from Jensen’s inequality and tower property of entropic risk measure which states that $\rho({{\boldsymbol{\zeta}}})=\rho(\rho({{\boldsymbol{\zeta}}}_{1}|{{\boldsymbol{\zeta}}}))$ for random variables ${{\boldsymbol{\zeta}}}_{1},{{\boldsymbol{\zeta}}}_{2}$. Note that this property is satisfied only by entropic risk measure in the family of law-invariant risk measures (Kupper and Schachermayer,, 2009). Notice that for large values of $K<N$, $N(1-\frac{1}{K})$ approaches $N$, thus the right-hand-side of (16) mimics the performance of the solution that uses all $N$ data points, that is, $\rho_{\mathbb{P}}(\ell({{\boldsymbol{z}}}^{*}(\hat{\mathbb{P}}_{N},\epsilon),{{\boldsymbol{\xi}}}))$.
@@ 5.1 Bias-aware cross validation
1:function K-foldCV($K,\mathcal{D}_{N},\epsilon$)
2:     $\mathcal{S}\leftarrow\emptyset$
3:     for $k\leftarrow 1$ to $K$ do
4:         $\mathcal{D}_{-k}\leftarrow\mathcal{D}_{N}\setminus\mathcal{D}_{k}$triangleright Training data (all samples except those in fold $k$)
5:         ^$\hat{{\mathbb{P}}}^{K}_{-k}\leftarrow$ empirical distribution of scenarios in $\mathcal{D}_{-k}$
6:         Solve problem (13) with distribution ^$\hat{{\mathbb{P}}}^{K}_{-k}$ and radius $\epsilon$ to get ${{\boldsymbol{z}}}^{*}(\hat{{\mathbb{P}}}^{K}_{-k},\epsilon)$
7:         $\mathcal{S}\leftarrow\mathcal{S}\cup\{\ell({{\boldsymbol{\xi}}},{{\boldsymbol{z}}}^{*}(\hat{{\mathbb{P}}}^{K}_{-k},\epsilon))\mid{{\boldsymbol{\xi}}}\in\mathcal{D}_{k}\}$
8:     end for
9:     return $\mathcal{S},\rho_{k\sim U(K)}(\rho_{{{\boldsymbol{\xi}}}\sim\hat{{\mathbb{P}}}_{k}^{K}}(\ell({{\boldsymbol{z}}}^{*}(\hat{{\mathbb{P}}}^{K}_{-k},\epsilon),{{\boldsymbol{\xi}}})))$
10:end function
To mitigate the underestimation of the entropic risk, we propose using the bias-aware bootstrap procedure described in Algorithm 1 together with either Algorithm 2 based on entropic risk matching, or with Algorithm 3 based on extreme value theory. Algorithm 5 describes our proposed approach for selecting the optimal $\epsilon$, with Algorithm 4 describing the $K$-fold CV step. One can recover the traditional biased CV procedure by setting $\delta=0$ in line 4 of Algorithm 5.
A pictorial representation of the debiasing effect in optimization problems can be seen in Figure 4. If the aim of Example 3 is to select the project with the lowest estimated risk, then most methods would favor project 3. However, by accurately estimating and correcting for bias–using our proposed approaches–project 1 becomes the preferred choice. Similar behavior is also observed when choosing $\epsilon$ using $K$-fold CV, where the problem parallels Example 3 in which the projects can be seen as corresponding to different regularization parameters $\epsilon$, with lower values of $\epsilon$ representing riskier projects. As we will see in the following section, the solution based on our proposed approach significantly outperforms traditional CV procedure.
@@
1:function RadiusTuning($\mathcal{D}_{N},K,M$)
2:     for $\epsilon\in\mathcal{E}$ do
3:         $\mathcal{S},\hat{\rho}\leftarrow\text{K-foldCV}(K,\mathcal{D}_{N},\epsilon)$
4:         $\delta\leftarrow\text{BootstrapBiasCorrection}(\mathcal{S},M)$triangleright Algorithm 1
5:         $\rho(\epsilon)\leftarrow\hat{\rho}+\delta$
6:     end for
7:     $\epsilon^{*}\leftarrow\arg\mathop{\rm min}_{\epsilon\in\mathcal{E}}\rho(\epsilon)$
8:end function
6 Distributionally robust insurance policy
The US National Flood Insurance Program (NFIP) provides flood coverage at subsidized premiums but faces significant challenges due to the large, correlated losses it insures against. These losses often result in claims exceeding the cumulative premiums collected over time (Marcoux and H Wagner,, 2023). Consequently, the NFIP currently operates with a deficit exceeding \$20 billion and is compelled to consider raising premiums (Marcoux and H Wagner,, 2023). However, higher premiums often deter households from purchasing coverage. This reluctance stems from how individuals perceive risk, which is frequently shaped by empirical losses rather than statistical estimates (Kousky and Cooke,, 2012). As a result, households tend to underestimate the risks associated with rare events. Demand for insurance, therefore, typically spikes only after catastrophic disasters (Gallagher,, 2014). In other words, individuals who have not experienced a catastrophic flood event are more likely to underestimate the associated risks. Surveys indicate that people exposed to flood risk but without firsthand experience of similar disasters often exhibit overly optimistic views about the threats posed by climate change. This optimism has been linked to houses in high flood-risk areas being overvalued by 6–9% (Bakkensen and Barrage,, 2022). Furthermore, NFIP premium subsidies reductions, combined with advances in risk estimation and flood risk mapping, have been shown to decrease house prices in high-risk areas (Hino and Burke,, 2021). Such behavioral responses are not unique to flood insurance markets. Herrnstadt and Sweeney, (2024) use a difference-in-differences method to show that the prices of houses within 500 feet of a gas pipeline in the Bay Area dropped by \$383 per household following the deadly 2010 pipeline explosion in San Francisco. Moreover, residents’ perceptions of risk increased significantly above the empirical average for several years after the explosion. Interestingly, however, the prices of properties located 2,000 feet from the pipeline remained unaffected, despite being classified as high-risk. This discrepancy underscores the impact of firsthand experiences of catastrophic events on risk perception.
We consider an insurance pricing problem, with one risk-averse insurer and $M$ representative risk-averse households. The proposed model can account for correlated losses and asymmetry in perception of risk measured by the empirical loss distributions at household level. Let $\alpha_{h}$ denote the risk aversion of household $h$, and let $\alpha_{0}$ represent the insurer’s risk aversion parameter. The uncertain loss faced by household $h$ is represented by $\xi_{h}$. The insurer offers a policy $(z_{h},\pi_{h})$ to household $h$, where the indemnity function $z_{h}\xi_{h}$ specifies the coverage provided for their loss $\xi_{h}$, and $\pi_{h}$ is the corresponding premium paid by household $h$. Consequently, the net loss faced by household $h$ under this policy is given by $\pi_{h}+(1-z_{h})\xi_{h}$. Let ^$\hat{{\mathbb{P}}}_{h,N}$ be the empirical distribution of losses faced by household $h$. The insurer’s demand response model assumes that household $h$ will accept the policy $(z_{h},\pi_{h})$ if the empirical entropic risk with insurance is less than the entropic risk without insurance. This condition is expressed by the following constraint:
$\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\pi_{h}+(1-z_{h})\xi_{h}\right)\leq\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\xi_{h}\right).$ (17)
The insurer aims to minimize risk across the policies offered to all households. Accordingly, the insurer’s true entropic risk is given by $\rho^{\alpha_{0}}_{\mathbb{P}}\left(\boldsymbol{z}^{\top}{{\boldsymbol{\xi}}}-{{\boldsymbol{1}}}^{\top}\boldsymbol{\pi}\right)$, where ${{\boldsymbol{1}}}$ is the vector of ones of the appropriate dimension. Since the true joint distribution, $\mathbb{P}$, of losses across all households with marginals $\mathbb{P}_{h}$, is unknown, the insurer replaces it with the empirical distribution ^$\hat{{\mathbb{P}}}_{N}$ and solves the following optimization problem to determine the policies offered to the $M$ households:
$\begin{array}[]{lll}\mathop{\rm min}&\rho^{\alpha_{0}}_{\hat{{\mathbb{P}}}_{N}}\left(\boldsymbol{z}^{\top}{{\boldsymbol{\xi}}}-{{\boldsymbol{1}}}^{\top}\boldsymbol{\pi}\right)\\ \text{s.t.}&\boldsymbol{\pi}\in\mathbb{R}_{+}^{M},\,\boldsymbol{z}\in[0,1]^{M}\\ &\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\pi_{h}+(1-z_{h})\xi_{h}\right)\leq\rho^{\alpha_{h}}_{\hat{{\mathbb{P}}}_{h,N}}\left(\xi_{h}\right)&\forall h\in[M].\end{array}$ (18)